OpenStudy (anonymous):

Use the Laws of Exponents to simplify the expression 5 to the 40th power over 5 to the 10th power Explain how you found your answer and write the answer as an exponent.

OpenStudy (anonymous):

since the base is the same (5 in this case), you subtract the exponents

OpenStudy (mertsj):

\[\frac{5^{40}}{5^{10}}\]

OpenStudy (anonymous):

5^40 / 5^10 = 5^(40 - 10)